Reports 1-1 of 1 Clear search Modify search
MIF (General)
tomotada.akutsu - 13:18 Tuesday 10 August 2021 (17865) Print this report
Shut the green beam X on the PSL table for investigating FNC

Follow-up of 17748. Maybe related to 17853.

Put a beam dump in front of the reflection mirror at the one arm’s edge of the FNC Michelson for X (see the attached photo). Let’s see.

By the way, the screw entangled with a black dog clamp in front of the reflection mirror seems kind of stuck… I tried to detach this because it did not work anymore for any purpose, but it was hard so I left it. This may be one of noise sources.

Images attached to this report
Comments to this report:
ryousuke.sugimoto - 14:07 Tuesday 10 August 2021 (17867) Print this report

About the clamp, I screwed it up three years ago when we installed ALS...
I’m so sorry.

takahiro.yamamoto - 13:29 Wednesday 11 August 2021 (17878) Print this report
Just after shutting out the beam from reference path, laser power fluctuation (K1:ALS-X_ARM_INPUT_IN1_DQ) was able to be seen on POP table.
As shown in Fig.1, it seemed to come from the change in the room temperature due to enter the PSL room.

After calming down the temperature fluctuation, laser power fluctuation were able to be seen on POP table (~10%) and PDA1 and PDA2 in PSL room (as a factor of 2~3).

When fiber path was shut out on 8/4, laser power fluctuation didn't seen on PDA1 and PDA2 in PSL room as shown in Fig.2. This fact mean that the laser power at the output of Prometheus is not so unstable.
(Around -17h in Fig.2, fiber path was shut out.)
So the fiber path (fiber coupling) seem to be the 1st candidate of causes of the amplitude fluctuation of Michelson fringes.
Images attached to this comment
kiwamu.izumi - 11:17 Tuesday 05 October 2021 (18451) Print this report

K. Izumi and S. Miyoki

The beam dump that Akutsu san had put in this past August (see the origianl post at 17865) was taken out today. We did it sometime in 10:00-11:00 in JST today. The main objective is to bring the fiber noise setup back to the nominal condifition so that we can further re-commission the system. Miyoki-san was in the PSL room at the same time but for other purpose.

Checking the Michelson fringe signals, we start seeing the interference fringe as expected. The attached is 10-day trend of the relevant signals showing the significant increase in the signals at the last minutes due to the removal of the dump today.

Images attached to this comment
Search Help